OpenStudy (anonymous):

Factor 8y2 + 17y + 2. (8y + 1)(y + 2) (2y + 1)(4y + 2) (2y + 2)(4y + 1)

OpenStudy (anonymous):

(8y +1)(y +2) because 2 * 8y = 16y... 1 * y = y... 16y + y = 17y and 1 * 2 gives 2

OpenStudy (anonymous):

Shannon, do you understand FOIL(First - Outer - Inner - Last)?

OpenStudy (anonymous):

a lil

OpenStudy (anonymous):

Well, you want to start with the first two terms in the factor components, so you should multiply which number by which number?

OpenStudy (anonymous):

okay